In topological space (N, T) where N is set of all natural numbers and
T {0, N, A, = {1,2,3, .., n}: n ∈ N}
Let A = {1, 2,4,6}, B = {5,7,9, 20}, find
A°, ext(A), b (A), B°, ext (B), b (B)
1
Expert's answer
2020-11-13T15:58:47-0500

"A^o" = union of open subsets of A. Hence "A^o= \{1,2\}\cup\{1\}\cup \{ \emptyset \}=\{1,2\}." ext(A)= union of open sets disjoint from A="\emptyset" since every nonempty set contains 1. b(A)= "\overline{A}\setminus A^o= \mathbb{N}\setminus \{1,2\}". since ext(A) is empty, closure of A is the entire space.

"B^o=" "\emptyset" since all non-empty open sets contain 1. ext(B)="=\emptyset\cup \{1\}\cup\{1,2\}\cup\{1,2,3\}\cup\{1,2,3,4\}" "=\{1,2,3,4\}." b(B)= "\overline{B}\setminus B^o= \overline{B}." Since "\{1,2,3,4\}"

is the largest open set disjoint from B. Hence "\overline {B}=\mathbb{N} \setminus\{1,2,3,4\}."